The body weights (kg) of 20 pregnant mothers attending to an antenatal Clinic were reported as Follows. 55,60,62,65,70,85,60,70,56,63,65,60,60 ,59,72,80,62,70,82,62 


1.Construct the class intervals with 7 classes.

2 Find the cumulative frequency

3 Create frequency distribution table

4 find mean, mode ,median

 






1
Expert's answer
2021-01-03T14:47:57-0500

1.

55, 56, 59, 60, 60, 60, 60, 62, 62, 62, 63, 65, 65, 70, 70, 70, 72, 80, 82, 85

Range of scores "=85-55 =30"

Number of classes "=7"

Class width "=\dfrac{30}{7}=5"

"\begin{matrix}
 \begin{alignedat}{2}
 Class \\
 limits
\end{alignedat} & \begin{alignedat}{2}
 Class\ \ \ \ \ \ \\
 boundaries
\end{alignedat} & Tally & Frequency \\
 53-57 & 52.5-57.5 & || & 2 \\
58-62 & 57.5-62.5 & \cancel{|||||}\ ||| & 8 \\
63-67 & 62.5-67.5 & ||| & 3 \\
68-72 & 67.5 - 72.5 & |||| & 4 \\
73-77 & 72.5-77.5 & & 0 \\
78-82 & 77.5-82.5 & || & 2 \\
83-87 & 82.5-87.5 & | & 1

\end{matrix}"


2.

"\begin{matrix}
 \begin{alignedat}{2}
 Class \\
 limits
\end{alignedat} &\begin{alignedat}{2}
 Frequency \\
 
\end{alignedat} & \begin{alignedat}{2}
 Cumulative \\
 frequency
\end{alignedat} \\
 53-57 & 2 & 2 \\
58-62 & 8 & 10 \\
63-67 & 3 & 13 \\
68-72 & 4 & 17\\
73-77 & 0 & 17\\
78-82 & 2 & 19 \\
83-87 & 1 & 20

\end{matrix}"


3.

"\begin{matrix}
 \begin{alignedat}{2}
 Class \\
 limits
\end{alignedat} & Frequency & \begin{alignedat}{2}
 Cumulative \\
 frequency
\end{alignedat} & \begin{alignedat}{2}
 Percentage \\
 
\end{alignedat} & \begin{alignedat}{2}
 Cumulative \\
 percentage
\end{alignedat} \\
 53-57 & 2 & 2 & 0.1 & 0.1 \\
58-62 & 8 & 10 & 0.4 & 0.5 \\
63-67 & 3 & 13 & 0.15 & 0.65 \\
68-72 & 4 &17 & 0.2 & 0.85 \\
73-77 & 0 & 17 & 0 & 0.85 \\
78-82 & 2 & 19 & 0.1 & 0.95 \\
83-87 & 1 & 20 & 0.05 & 1

\end{matrix}"


4.


"mean=\dfrac{1}{20}(55+56+59+4(60)+3(62)+63"

"+2(65)+3(70)+72+80+82+85)=65.9"


"mode=60, \text{appeared 4 times}"





"median=\dfrac{62+63}{2}=62.5"